# Low-depth Clifford circuits approximately solve MaxCut

### v2

@@ VI.2 Performance on unweighted Erdös-Rényi graphs
First, we fixed the edge probability to $1/2$ and study the performance with respect to the problem size. In Fig. 5e we show the instance averaged minimum energy of solutions obtained with randomized ADAPT-Clifford (green) and standard GW (light grey). For graphs up to $N=1000$. The randomized version of our algorithm produces better solutions, on average, than GW, an observation that is verified by the instance-averaged minimum energy differences shown in Fig. 5h.
@@ VII Discussion and outlook
We introduce ADAPT-Clifford, a quantum inspired classical approximation algorithm for MaxCut. For each problem instance, ADAPT-Clifford builds a low-depth Clifford circuit to prepare a stabilizer state that encodes an approximate solution. The algorithm was inspired by observation of the (almost) Clifford character of the ADAPT-QAOA solution circuits for MaxCut on weighted fully connected graphs. We introduce a randomized and a deterministic variant of this algorithm. Their respective runtime complexities are $O(N^{2})$ and $O(N^{3})$ for sparse graphs, and $O(N^{3})$ and $O(N^{4})$ for dense graphs, and in all cases the space complexity is $O(N^{2})$. Naturally, the deterministic variant always outperforms the randomized variant, albeit at the cost of an increased runtime.
@@
We hope the results reported here will help delimit the subset of graphs where a quantum speedup could be expected and thus where the current efforts should focus, in similar spirit to previous results obtained with a different subuniversal family of gates [111]. While our work indicates that ADAPT-Clifford has a guaranteed approximation ratio, we do not yet have a proof. Whether it is possible to improve the runtime without compromising the solution quality is also an open problem. Our algorithm showed the poorest performance on fully connected graphs with signed weights. This was anticipated in Sec. III since the ADAPT-QAOA solution circuits are not fully Clifford. However, they are near-Clifford, motivating then a resource-centered design of variational ansätze, with a Clifford mixer part constructed following a scheme like the one introduced in this work, similar in spirit to the optimal mixers restricted to subspaces [112], and a cost part with few variational parameters adding just the right amount of nonCliffordness necessary to approximate the problem up to a desired ratio. Furthermore our algorithm could aid in reducing the cost of parameter optimization in QAOA when used to warm-start [113] it. Finally, our Clifford algorithm was tailored to solve the MaxCut problem. It remains an open question to what extent other combinatorial optimization problems admit Clifford approximation algorithms with practical polynomial runtimes.
@@ Appendix A The Goemans-Williamson algorithm
All the results for the GW algorithm reported in this manuscript have been obtained using a freely available Julia implementation  [^10] It can be found at https://github.com/ericproffitt/maxcut.jl.
@@ Appendix C Gradient computation without explicit evaluation of the expectation values
We would like to point out that the explicit form of the gradient in Eq. (31) establishes a connection with the a family of heuristic algorithms for MaxCut known by the name of Sahni-Gonzalez [87, 88]. This connection was recently studied in Ref. [89].
@@
Appendix D Some explicit examples of ADAPT-Clifford solving MaxCut
@@
D.1 An example with $N=5$ nodes
@@
D.2 An example with $N=4$ nodes
@@
Appendix E Estimation of the mean approximation ratios for the case of positive weights

### v3

@@ VI.2 Performance on unweighted Erdös-Rényi graphs
First, we fixed the edge probability to $1/2$ and study the performance with respect to the problem size. In Fig. 5d we show the instance averaged minimum energy of solutions obtained with randomized ADAPT-Clifford (green) and standard GW (light grey). For graphs up to $N=1000$. The randomized version of our algorithm produces better solutions, on average, than GW, an observation that is verified by the instance-averaged minimum energy differences shown in Fig. 5h.
@@ VII Discussion and outlook
We introduce ADAPT-Clifford, a quantum inspired classical approximation algorithm for MaxCut. For each problem instance, ADAPT-Clifford builds a low-depth Clifford circuit to prepare a stabilizer state that encodes an approximate solution. The algorithm was inspired by observation of the (almost) Clifford character of the ADAPT-QAOA solution circuits for MaxCut on weighted fully connected graphs. A comparison between the two methods and resource estimates for ADAPT-Clifford are given in App. E. We introduce a randomized and a deterministic variant of this algorithm. Their respective runtime complexities are $O(N^{2})$ and $O(N^{3})$ for sparse graphs, and $O(N^{3})$ and $O(N^{4})$ for dense graphs, and in all cases the space complexity is $O(N^{2})$. Naturally, the deterministic variant always outperforms the randomized variant, albeit at the cost of an increased runtime.
@@
We hope the results reported here will help delimit the subset of graphs where a quantum speedup could be expected and thus where the current efforts should focus, in similar spirit to previous results obtained with a different subuniversal family of gates [110]. While our work indicates that ADAPT-Clifford has a guaranteed approximation ratio, we do not yet have a proof. Our algorithm showed the poorest performance on fully connected graphs with signed weights. This was anticipated in Sec. III since the ADAPT-QAOA solution circuits are not fully Clifford. However, they are near-Clifford, motivating then a resource-centered design of variational ansätze, with a Clifford mixer part constructed following a scheme like the one introduced in this work, similar in spirit to the optimal mixers restricted to subspaces [111], and a cost part with few variational parameters adding just the right amount of nonCliffordness necessary to approximate the problem up to a desired ratio. Furthermore our algorithm could aid in reducing the cost of parameter optimization in QAOA when used to warm-start [112] it. More broadly, Clifford circuits can be leverage to construct a framework for the efficient state initialization in variational quantum algorithms beyond the product state paradigm. An example of this applied to quantum chemestry problems was introduced in Ref. [113]. Finally, our Clifford algorithm was tailored to solve the MaxCut problem. It remains an open question to what extent other combinatorial optimization problems admit Clifford approximation algorithms with practical polynomial runtimes.
@@ Appendix A The Goemans-Williamson algorithm
All the results for the GW algorithm reported in this manuscript have been obtained using a freely available Julia implementation [114].
@@ Appendix C Gradient computation without explicit evaluation of the expectation values
We would like to point out that the explicit form of the gradient in Eq. (31) establishes a connection with a family of heuristic algorithms for MaxCut known by the name of Sahni-Gonzalez [87, 88]. This connection was recently studied in Ref. [89].
Appendix D Time to solution analysis: ADAPT-Clifford vs Goemans-Williamson
Figure 9: (a) Problem instance averaged time to solution for randomized ADAPT-Clifford (orange solid line) and GW with $\mathcal{I}=10^{0},10^{5}$ (dashed light grey and black), for the weighted complete graphs considered in Fig. 2. (b) Time to solution of GW as a function of the number of times the rounding step $I$ is performed. Results are shown for three different system sizes, $N=300,400,500$ from bottom to top, respectively. The empty circles correspond to all the individual TTS for the different $100$ problem instances solved per problem size, the full circle to the mean TTS and the error bars shown the standard deviation. The dashed lines are guides for the eye.
In order to complete the comparison between the performance of ADAPT-Clifford and GW, we conducted a time to solution (TTS) analysis. This study was carried out on an Apple M1 Pro laptop with 8 cores. We focused on the case of weighted complete graphs studied in Sec. V.1, and compared our Python based implementation of ADAPT-Clifford available at [90] against the freely available Julia implementation of GW for the MaxCut problem [114]. Since the TTS of deterministic ADAPT-Clifford can be easily determined as one order of magnitude larger than that of randomized ADAPT-Clifford, we restricted the TTS study to the latter variant of our algorithm. For the GW we restricted ourselves to consider the situations of only $\mathcal{I}=10^{0}$ and $\mathcal{I}=10^{5}$, the two extreme values taken for the results in Sec. V.1.
The results are shown in Fig. 9. Several observations are in order. The TTS of GW shows a relative slow increase with system size for small problems, until about $N\sim 300$ after which the familiar $O(N^{3})$ scaling is observed. Additionally, we do not find a strong dependence of the TTS with $\mathcal{I}$, in fact for the problem sizes considered in this work, i.e., up to $N=1000$, the TTS does not increase with $\mathcal{I}$ long as this is not bigger than $O(10^{2})$, see Fig. 9b. When the value of $I$ exceeds this threshold, we do observe an increase in the runtime albeit not a considerable one. As such, deterministic ADAPT-Clifford will only be superior both in solution quality and TTS up to $N=30$.
For randomized ADAPT-Clifford the TTS is always better than GW, see Fig. 9a. In fact, we empirically find a more advantageous scaling of $O(N^{2.7})$ when contrasted to the $O(N^{3})$ found in the same manner for GW. However, this makes the comparison considerably more subtle. Although the quality of solution found by GW is already better than randomized ADAPT-Clifford when $\mathcal{I}>10$, GW relies on a Cholesky decomposition which has highly optimized numerical subroutines that exploit the multi-core nature of modern CPUs. However, randomized ADAPT-Clifford is based on STIM and thus runs on a single core. One could then easily improve the quality of solution without sacrificing the TTS or runtime scaling by executing the algorithm for different initial position in parallel. We thus believe, this variant of the algorithm does offer an advantage over GW.
Appendix E Resource estimates for ADAPT-Clifford
As mentioned in Sec. VII one might wonder how ADAPT-Clifford compares with ADAPT-QAOA? And what would be the resource demands to implement ADAPT-Clifford on a near-term device? In this appendix we explore these two questions.
Given the computational cost of directly simulating ADAPT-QAOA we restrict the comparison to small problem instances, as for example those solved in Fig. 1. As we observed in Sec. III, the best performing ADAPT-QAOA solution circuits are very close to Clifford. In this sense, the performance, quantified by the approximation ratio, of both ADAPT-QAOA and ADAPT-Clifford, for these small problem instances, must be the same. Now, if the comparison is extended other metrics, for instance number of layers or steps, cost of parameter optimization, cost of gradient evaluation, then ADAPT-Clifford will be best.
In order to see this we will show that ADAPT-Clifford can be directly obtained from ADAPT-QAOA by systematically reducing the ansatz expressivity. We start from the full ADAPT-QAOA ansatz in Eq. (7), and use the observations summarized at the end of Sec. III to systematically reduced the ansatz expressivity. First, we start from the state $Z_{k}\mathrm{H}^{\otimes N}|0\rangle^{\otimes N}$ where $k$ is chosen randomly. Second, we set all $\gamma_{l}=0$ and the $\beta_{l}=-\pi/4$, this guarantees the resulting circuit to be Clifford. Third, we restrict the operator pool in Eq. (9) to $\mathrm{P}_{\rm OP}=\{Y_{l}Z_{m},Z_{l}Y_{m}\}_{j,k=1,...,N,j\neq k},$, where one of the two indices is fixed by our choice of $k$. Fourth, we use the gradient Criteria to select the mixer Hamiltonian at any intermediate step.
The above algorithm is a direct simplification of the full ADAPT-QAOA ansatz, and although it is different from the algorithm presented in Sec. IV, it already allows us to see why ADAPT-Clifford is a more efficient algorithm. First, the gradient evaluation is highly simplified by the reduction of the operator pool size. Second, the need for parameter optimization is bypassed. Third, the restriction to Clifford unitaries enables efficient classical simulation. Finally, for ADAPT-Clifford we showed in App. C that the gradient evaluation can be done without the need for explicit evaluation of expectation values, which further speeds up the algorithm.
Let us now focus on the second question. We based the resource estimation on the CNOT count. First, let us consider a qubit chip with all-to-all connectivity. In this case, and given the gate shown in Eq. (21), it is easy to see that the total number of CNOT gates required to run the circuit is $2N$, with $N$ the number of nodes in the problem graph.
Consider now the opposite case, a qubit architecture with only linear connectivity. In order to apply a CNOT between two qubits at arbitrary positions, say $l$ and $m$, we bring the second qubit $m$ to $l+1$ using a swap network, then apply the Clifford gate, and swap the second qubit back to its original position $m$. In the following we use the fact that a swap can be implemented with $3$ CNOTs. Further, we assume that the two reference qubits, $k$ and $j$ are on opposite ends of the chain, thus $|k-j|=N-1$, and implementing the entangling Clifford gate between these two qubits requires $26(N-2)+$ 2 CNOT gates. Finally, we assume the worst scenario in terms of separation between the qubits, where the entangling Clifford gates need to be applied between the reference qubit and the qubit which is furthest from it, among those inactive. Thus, the rest of the $N-2$ steps can be implemented using $2(N-2)+6\sum_{s=0}^{N-3}s=3(N-3)(N-2)+2(N-2)$ CNOT gates. Which leads to a total CNOT count of $2+(N-2)\left[3(N-3)+8\right]$.
@@ Appendix E Resource estimates for ADAPT-Clifford
Appendix F Some explicit examples of ADAPT-Clifford solving MaxCut
@@
F.1 An example with $N=5$ nodes
@@
F.2 An example with $N=4$ nodes
@@
Appendix G Estimation of the mean approximation ratios for the case of positive weights